Forty percent of banks did not earn profits during 2003. The SEC selects two banks at random. Find the probability that neither of them earned profits in 2003.
Answer by stanbon(75887)

Forty percent of banks did not earn profits during 2003. The SEC selects two banks at random. Find the probability that neither of them earned profits in 2003.
----
Binomial Problem with n = 2 and p(no profit) = 0.4
---
P(x = 2) = 2C2*0.4^2*0.6^0 = 0.4^2 = 0.16
============
Cheers,
Stan H.
===========